OpenStudy (anonymous):

Simplify the expression. (3 – i) – (2 + 6i) (1 point) A.1 – 7i B.–1 + 7i C.–6i D.5 + 5i

OpenStudy (anonymous):

@jdoe0001 @agent0smith @DonaldRoyMiller @ybarrap

OpenStudy (ybarrap):

$$ (3 – i) – (2 + 6i) \\ =(3-2)+(-i+6i)\\ $$ I've separated the real and imaginary parts for you, just add each independently and you are done. For example, 3-2=1.

OpenStudy (anonymous):

oh ok i was like hold uppp hahah (: is it A?

OpenStudy (agent0smith):

Distribute the negative then combine like terms (3 – i) – (2 + 6i) 3 - i - 2 - 6i then you can change the order as ybarrap did,

OpenStudy (ybarrap):

$$ (3 – i) – (2 + 6i) \\ =(3-2)+(-i-6i)\\ $$

OpenStudy (agent0smith):

A is correct.

OpenStudy (anonymous):

Hi emma97

OpenStudy (anonymous):

ok thanks guys! @ybarrap @agent0smith heyyy! @DonaldRoyMiller

OpenStudy (anonymous):

Don't let them do it for you. Let's go through it. It'll just take a moment.

OpenStudy (anonymous):

hahah they didnt do it for me they just helped me out with realizing the steps i need to do! i eventually did the math lol!

OpenStudy (anonymous):

You can get where you can do it in your head. For instance, the (3 – i) isn't a problem. The problem lies in realizing you have to distribute that minus one. I'm pretty certain that was the point of the exercise--to get you to realize that. :)

OpenStudy (agent0smith):

@DonaldRoyMiller nobody did this for her. I showed ONE step.

OpenStudy (anonymous):

You DID do it for her. You could have pointed out the need to distribute, but you didn't.

OpenStudy (anonymous):

The problem is only two steps.

OpenStudy (agent0smith):

...k. Here is what i wrote: Distribute the negative then combine like terms (3 – i) – (2 + 6i) 3 - i - 2 - 6i

OpenStudy (anonymous):

Okay. I still think it's far better to ask questions, like what do you do with -(2+6i)? I think the whole point of the exercise was to get the student to realize they have to distribute that minus one. But, okay. Fair enough. I guess you didn't do it for her, exactly.

OpenStudy (agent0smith):

lol geez, saying "distribute the negative" is pretty clear imo.

OpenStudy (anonymous):

That's because you already know it?

OpenStudy (anonymous):

Anything is easy if you know it.

OpenStudy (agent0smith):

Hence why i showed what "distributing the negative" does: (3 – i) – (2 + 6i) 3 - i - 2 - 6i Look at some of my other answers - I ask questions... when i think it'll help the student get started. Some students just need to see that first step, then they're fine from there, and they'll recognize that first step next time.

OpenStudy (anonymous):

I think Socrates had it right. That's about all I can say.

OpenStudy (agent0smith):

And btw, "distribute the negative" is clear to almost all students.

OpenStudy (anonymous):

One, or both, of us is talking to a brick wall. Good day.